OpenStudy (anonymous):

kRachael went to the internet cafe traveling 6 mph and returned home traveling 4 mph. If the total trip took 5 hours, how long did Rachael travel at each speed? hour at 6 mph hour at 4 mph

OpenStudy (anonymous):

let S be the distance travelled therefore : \[\frac{S}{6}+\frac{S}{4} = 5 \] now simplify for S u will get the distance.

OpenStudy (anonymous):

the distance is 12.. now what formula should i use to find the time for each speed?

OpenStudy (anonymous):

velocity = distance/time

OpenStudy (anonymous):

t2 = 5 - t1 6 t1 = 4 ( 5 - t1) 10 t1 = 20 t1 = 2 hrs =>t2 = ...